Tuesday, November 29, 2005

Week of 11-21-05

Specific project tasks completed this week include:

- Happy Thanksgiving! Most project work done at home this week in between eating Turkey.

- Over the break, the original literature reviewed was looked at again as the project is at a point where significant changes are going to be made to the current working setup over the next month or so. Reviewing the previous work will help to help be sure that proposed changes (sinusoidal modulation using LCD and new laser) made will work properly.

- Follow up email with ColorLink regarding LCD display.

Week of 11-14-06

Specific project tasks completed this week include:

- Met with Dr. Blair regarding current project status. Based on measurements taken on the current optical system it was determined that the best course of action at this point will be to pursue intengrating sinsoidal amplitude modulation within the system. Investigation into electro-optical modulators proved them to be a price prohibitive option. It was suggested that liquid crystal displays could be used as transmission filters. Vendors of such displays include:

Boulder Nonlinear Systems
ColorLink

- Feasability of integrating such displays was explored through product information and vendors were contacted questioning display absorbtion at the wavelengths used in the slow light system.

- Literature previously reviewed was explored again to determine if a new laser which Dr. Blair could obtain operating at 473nm with a maximum 200mW operating power could be integrated into the experiement.

Week of 11-7-05

Specific project tasks complete this week include:

- Brief family vacation

- Based on previous measurements of propagation time shift (slow light) as well as newly obtained output power stability measurements, it can be seen that to create a robust system sinusoidal output modulation is very necessary. While square wave chopping of the output beam works reasonably well, in inherently introduces noise into the measurements and does not provide a set or extremely accurate way of measuring propagation shift. Electro-optical modulation or a similar means is a neccessity at this point in the experiment to make measurements more accurate and easier to take. EO modulators were investigated and several vendors were briefly looked at, including: Electro-Optical Products Corporation, New Focus, Leysop, and Global Spec.

- It looks as if these modulators may be very price prohibitive!

Week of 10-31-05

Happy Halloween!

Specific project tasks completed this week include:

- Measurements were taken using the new high sensitivty detectors obtained the previous week. Power stability of the system was measured at the same time scale (every 1 min for 1 hr) as the previously taken measurements, and two of such measurements were taken. It was observed that output power stability for the first 20 - 25 minutes fluctuated somewhat, but these effects were minimized after approxmately 30 minutes of warm up time.

- Based on this information, a standard warm up time of 30 minutes for the system will be implemented during all future experiments.

Tuesday, November 01, 2005

Week of 10-24-05

Specific project tasks completed this week include:

- Measurement data on laser output power stability taken last week was analyzed. It was found that any output power instability that could be the cause of the propagation time shift noise jitter through the crystal that has been observed would be fairly minimal, and below the detection sensitivity of the detectors used last week (ThorLabs). Higher sensitivity detectors were located and new measurements on power stability will be taken (same time scale - every 1 min for 1 hr).

- Cleaned up the lab!

- Midterms.........